# Cost-driven prunings for iterative solving of constrained routing problem with SRLG-disjoint protection

### v1

@@ 28            $c_{up}\ +=\ cost\_step$
First of all, note, that the problem statement for SRLG-disjoint DRCR problem does not require PP to have optimal cost. So, there is no need to use the full version of Pulse+ in the search for PP. Instead we propose a simplified version of Pulse+, that returns the first found feasible protection path. It still contains the infeasibility pruning by delay and guarantees the satisfaction of constraints, but it does not track the cost of the found path and does not update the found solution to minimize the cost. More details of it can be found in the Appendix A-C.
@@ IV-A Dataset and metrics
We evaluate performance of our algorithms, using dataset published by authors of paper [22]. It is available by the link [27] (link in the original paper is not available). The dataset contains network topologies of two types: real-world and randomly generated networks. Real-world networks originate from the Topology Zoo collection [28]. Random networks are Erdos-Renyi (ER) graphs with different number of nodes and average densities. There are seven real-world instances from the Zoo, all relatively small, with the biggest having only 754 nodes. Random networks are generated with 1000, 2000, 4000, 6000, 8000, and 10000 nodes and three different average density values. This dataset is further duplicated and adapted separately for the DRCR problem and for SRLG-disjoint DCR problem. For the latter, two different generators of SRLG are applied, following ”random” and ”star” patterns [29], [30].
@@
Name Nodes Links
Cogentco 197 490
GtsCe 149 386
Interoute 110 316
Kdl 754 1798
Pern 127 258
TataNld 145 388
VtlWavenet2008 88 184
@@
For the ”random” generator, the size of an SRLG is a random integer between 1 and 40, and links are randomly chosen across the whole graph to be included in the SRLG. The procedure is repeated until each link in the graph is included in at least one SRLG. For the ”star” generator, links to include in an SRLG are chosen randomly from egress (outgoing) links of a particular node in the graph, and the size of an SRLG is randomly determined based on the average degree of the graph. As mentioned in [22], the ”star” generator represents optical networks, while random may originate from other physical backgrounds.
@@
There are 10 independent instances of random graphs for each combination of node number and density in DRCR case, making total of 180 graphs (10 random instances for each one of 6 node number options combined with 3 density options), and 3 independent instances for each set of parameters in SRLG case, making total of 108 graphs (3 random instances for each one of 6 node number options combined with 3 density options and 2 SRLG generators).
@@
We extend the dataset, adding another important example of random graphs, namely Barabasi-Albert graphs. Those graphs show scale-free behavior [31], therefore we call this extended dataset scale-free (SF). Since the work of Barabasi and Albert, there has been a broad discussion on the extent to which such graphs reflect the properties of real-world networks. Without doing a complete review, we can mention works [32], [33], [34] that can give some insights into the discussion. Here, we make a simple argument that this class of networks is important and serves as a good complement to the two classes of networks present in the original dataset. For a comparison of the three types of networks, see Appendix B. We used the Python NetworkX library to generate SF networks and followed the general design established in the original dataset. We also chose the number of nodes to be 1000, 2000, 4000, 6000, 8000, and 10000. For each number of nodes, we chose parameter m of the generator to be 2, 3, and 4 and generated 10 independent instances for each option. Then, we turned the graphs to be directed by adding an edge in the opposite direction for each edge in the generated graph. Finally, we assigned costs and delays to be independent uniformly distributed integers in the range $[1,...100]$. For SRLG problems, we chose the first five graphs in each category and independently generated random and star SRLGs. Then, we generated 50 source-target pairs for each graph and assigned delay constraints for them. Finally, for the DRCR problem, we removed all tasks that appeared to have no feasible solution, and for SRLG-disjoint DCR problem, we kept only trap instances. After that, the total number of DRCR instances is 8498, the total number of trap instances for ”random” SRLG is 3404, and for ”star” SRLG is 3264. We consider non-trap instances trivial and thus not representative for evaluating the BTCS algorithm.
@@ IV-B DRCR
All instances of the DRCR problem in the dataset from [22] are relatively easy to solve for the Pulse+ algorithm. So, we do not split it into parts according to different topology types but just generate statistics for the entire dataset altogether. This dataset has a total number of 11245 tasks, and the results are presented in Table III.
@@
We run four algorithms on this dataset. The first algorithm in the table is the Pulse+, implemented by the authors of [22] and obtained on Gitee [27]. Second is the Pulse+, which we developed independently. Surprisingly, these two significantly differ in performance, which we assume to be an implementation flaw since all main features of our Pulse+ implementation follow the ideas of [22]. Our code is built using Boost Graph Library, while the code in [27] uses its own data structure for graph representation.
@@ IV-C SRLG-disjoint DCR problem
To highlight algorithm performance when dealing with different network topologies and different SRLG assignment options, we present the results separately for real-world, ER, and SF networks, each with random and star SRLG generators. There are six combinations, and the results are listed in tables V - IX. The only combination missing in these tables is real-world networks with random SRLG, which turned out to have no feasible solutions. To show the advantage of parallel execution of BTCS and give further insight into its nature, we run it in 1, 2 and 4 parallel threads. For the corridor width in all experiments we use the cost of the shortest edge in the graph, multiplied by 10.
@@ IV-D Analysis and discussion
IV-D1 Single-thread performance
All results in the previous section are presented for single- and multithread implementation of the BTCS algorithm. Importantly, considering the number of feasible solutions found under the limit of 20 or 50 ms, BTCS algorithm in most cases has advantage over CoSE-Pulse+ even in single-thread version. The only exception is ER instances with random SRLG, where the single-thread performance is inferior (though still comparable). Considering other metrics, such as maximal, average and median time to feasible solution, BTCS in single thread also has comparable performance to CoSE-Pulse+, and outperforms it in certain scenarios (see Tables V - IX for details). The only exception is the SF instances with star SRLG, where the maximal execution time of BTCS in single thread is almost 6x worse than that of CoSE-Pulse+. This observation underpins the claim that the competitiveness of BTCS originates from its basic principles and not from any sophisticated multithread implementation.
IV-D2 Multithread performance
The multithread execution results in the previous section are obtained with a simple and relatively straightforward approach. In this implementation, the algorithm performs parallel search for AP candidates in several subsequent cost corridors. Other variants, such as the parallel search for disjoint PP within a corridor, are possible, as mentioned in Section III-D. CoSE paradigm also allows parallelization, as in [21], although it was not generalized to handle delay-constrained problems to the best of our knowledge. We consider detailed research of parallel execution capabilities for the SRLG-disjoint DRCR problem a subject of future work.
The approach implemented in the present work allows further insight into how the distribution of feasible solutions influences the algorithm’s performance. Note that Tables V - VII show ”normal” behavior, i. e. the number of found feasible solutions increases and execution time decreases with growing number of threads. On the contrary, Tables IX - IX show ”abnormal” behavior, namely, the performance metrics deteriorate between two and four threads. We attribute this issue to the path distribution relative to cost. Too many threads means that the higher thread will often search in the space above the shortest feasible solution. Since there are more AP candidates at higher cost values, these threads will take more time, slowing down the whole process.
To understand this phenomenon more deeply, we introduce parameter $\alpha$ to account for the corridor width in the algorithm implementation. The width of the corridor is the cost of the shortest edge in the graph, multiplied by $\alpha$. So, the experiments performed in the previous section correspond to $\alpha=10$. In Figure 5, we plot the number of feasible paths found under 20 and 50 milliseconds as a function of this parameter $\alpha$, which is effectively the corridor width. The figure shows that algorithm performance in one and two threads is quite stable with respect to $\alpha$ and changes only a little in the wide range of parameter values. The value $\alpha=10$ is close to optimal for these two options, and double-thread implementation has clear advantage over single-thread. At the same time, the four-thread implementation does not have an overall advantage compared to double-thread, and its performance decays quickly with growing $\alpha$.
@@ IV-D3 Worst-case performance
Figure 6: Distribution of solution times for SRLG-disjoint DCR problem on the real-world network instances. (a): absolute execution time of BTCS algorithm on Zoo dataset with star SRLG. (b): distribution of execution times for CoSE-Pulse+ algorithm against single-thread BTCS. (c): BTCS algorithm execution time dependency on the graph size. (d): BTCS algorithm execution time dependency on the number of cost corridors explored before the solution is found. (c): BTCS algorithm execution time dependency on the number of AP candidates that are shorter than the solution, but have no disjoint protection.
Although worst-case performance is interesting theoretically, it is a less valuable metric in real applications. For an algorithm intended to be applied in the operation of modern networks, with acceptable delays measured in milliseconds, an algorithm that solves a problem in a second does not gain much advantage over the one that takes 5 seconds for the same task. This is why we decided to primarily compare algorithms regarding the number of problems solved in 20 and 50 milliseconds. Nonetheless, we provide and discuss some data below to give insight into the algorithm’s worst-case performance.
@@ V Conclusion and future work
We suggest several ways of future improvement. One of them is application of the technique inspired by the queuing procedure from paper [15]. The original procedure stops pulses that have reached certain number of links, placing them in a queue and reusing again after all pulses have reached the limit. This approach turns the DFS search in Pulse algorithm into a combination of DFS and BFS, making it pure BFS when the depth limit is set to 1. This can be used as an inspiration in the search for AP candidates in BTCS, by placing the pulses that have reached the upper limit in cost for current corridor in a queue and reusing in the next corridor instead of pruning them. It will improve speed at the price of memory to store all the pulses, that otherwise would be pruned.
Also it may be possible to use observations on search space structure made in this work to speed up the CoSE algorithm. A merging of these two techniques, exploiting conflicting link sets together with judiciously chosen cost bounds for efficient pruning, may result in further progress in the area.
Another way of improvement, considering the multithread implementation of the algorithm, is the development of more sophisticated ways of interaction between threads that will allow the stopping of execution at the moment when the solution is found. It will also allow mitigation of the performance decline for higher number of threads, observed on SF networks. Carefully designing breaking rules for multithread execution, one can also turn the algorithm into an approximate one, trading exactness of the solution for speed. For example, it can be achieved by independent exploration of odd and even corridors in two parallel threads, allowing one thread not to wait for the other on each step and breaking at the time when the first protected pair is found. Since the exploration goes in the bottom-top direction, the solution found will be relatively close to the optimal.
Finally, we can think of including parallelism on different level, and instead of searching for candidates in parallel corridors, search for protection paths in parallel for all AP candidates in each corridor, as was discussed in Section III-B.
@@ A-A Pulse+ algorithm
The pseudocode, presented here, corresponds to the original Pulse+ algorithm, proposed in paper [22]. It is intended to solve Delay-Range Constrained Routing problem and return the optimal path for it. In our paper it is used twice: in section III-B (Algorithms 1 and 2) and in Algorithm 3. We divide the code into two parts for convenience: the high-level code (Algorithm 4) and details of recursive Pulse function (Algorithm 5).
@@ 5      $current\_min\_cost=c$
Lines 5 and 5 of algorithm 5 contain infeasibility and optimality prunings, respectively. Dominance pruning is not applicable to the scenario with nonzero $d_{low}$, so, generally this algorithm is not as efficient as original Pulse. In order to avoid loops in found paths, it is also essential to track visited nodes in this scenario.
@@ A-B Pulse+ modification for all paths search in cost corridor
On the line 3 of algorithm 3 we use modification of Pulse+ algorithm to save all paths in current cost corridor. We put it into Appendix, because the difference with original Pulse+ is minor.
@@ A-C Pulse+ algorithm for first feasible path
In this section we describe, perhaps, the simplest implementation of the Pulse+ algorithm, used in algorithm 3 on line 3. It is intended to return any feasible path as fast as possible. So, it does not keep track of cost value, and applies only infeasibility pruning. It also must terminate the recursion, when the first feasible path was found. The function on line 9 of algorithm 9 is responsible for it, but we do not go into details of implementation. On practice it can be done by inclusion of a global boolean flag or by any other mean.

### v2

@@ 27            $c_{up}\ +=\ cost\_step$
First of all, note, that the problem statement for min-min SRLG-disjoint DRCR problem does not require PP to have optimal cost. So, there is no need to use the full version of Pulse+ in the search for PP. Instead we propose a simplified version of Pulse+, that returns the first found feasible protection path. It still performs the infeasibility pruning by delay and guarantees the satisfaction of constraints, but it does not track the cost of the found path and does not update the found solution to minimize the cost. More details of it can be found in the Appendix A-C.
@@ IV-A Dataset and metrics
We evaluate performance of our algorithms, using dataset that contains networks published with the paper [22] (check the link [30]), and a set of scale-free networks generated by ourselves.
The dataset [30] contains network topologies of two types: real-world and randomly generated networks. Real-world networks originate from the Topology Zoo collection [31]. Random networks are Erdos-Renyi (ER) graphs with different number of nodes and average densities. There are seven real-world instances from the Zoo, all relatively small, with the biggest having only 754 nodes. Random networks are generated with 1000, 2000, 4000, 6000, 8000, and 10000 nodes and three different average density values. This dataset is further duplicated and adapted separately for the DRCR problem and for SRLG-disjoint DCR problem. For the latter, two different generators of SRLG are applied, following ”random” and ”star” patterns.
@@
Name Nodes Links Planar
Cogentco 197 490 Not
GtsCe 149 386 Yes
Interoute 110 316 Yes
Kdl 754 1798 Not
Pern 127 258 Yes
TataNld 145 388 Yes
VtlWavenet2008 88 184 Yes
@@
”Star” SRLG pattern, also called ”incident-SRLG”, is widely known in literature on optical network survivability [21], [32], [33], [34]. For the ”star” generator, links to include in an SRLG are chosen randomly from egress (outgoing) links of a particular node in the graph, and the size of an SRLG is randomly determined based on the average degree of the graph. For the ”random” generator, the size of an SRLG is a random integer between 1 and 40, and links are randomly chosen across the whole graph to be included in the SRLG. The procedure is repeated until each link in the graph is included in at least one SRLG. Although we do not know literature except for the work [22] to explicitly study random SRLG, we believe that this rather simple technique may be an example of ”non-star” pattern mentioned in work [33], and it provides a complementary example to the ”star” pattern to improve diversity of instances in the dataset.
We should also mention another widely-known SRLG pattern, namely ”regional” SRLGs [28], [29], [35], [36]. Although being important for certain practical scenarios, it is mainly applicable to planar or nearly-planar network topologies and requires additional geographical information on the network node location. It is not present in dataset [30] and is not used in our work.
@@
The dataset [30] contains 10 independent instances of random graphs for each combination of node number and density in DRCR case, making total of 180 graphs (10 random instances for each one of 6 node number options combined with 3 density options), and 3 independent instances for each set of parameters in SRLG case, making total of 108 graphs (3 random instances for each one of 6 node number options combined with 3 density options and 2 SRLG generators).
@@
We extend the dataset, adding another important example of random networks, namely Barabasi-Albert graphs. Those graphs show scale-free behavior [37], therefore we call them scale-free (SF). Since the work of Barabasi and Albert, there has been a broad discussion on the extent to which such graphs reflect the properties of real-world networks. Not doing a complete review, we mention works [38], [39], [40] that give some insights into the discussion. We make a simple argument that this class of networks is important and serves as a good complement to the two classes of networks present in the original dataset. For a comparison of the three types of networks, see Appendix B. We used the Python NetworkX library to generate SF networks [41] and followed the general design established in [22]. We also chose the number of nodes to be 1000, 2000, 4000, 6000, 8000, and 10000. For each number of nodes, we chose parameter m of the generator to be 2, 3, and 4 and generated 10 independent instances for each option. We turned the graphs to be directed by adding an edge in the opposite direction for each edge in the generated graph. Costs and delays are set to be independent uniformly distributed integers in the range $[1,...100]$. For SRLG problems, we chose the first five graphs in each category and independently generated random and star SRLGs. We generated 50 source-target pairs for each graph and assigned delay constraints for them. Finally, for the DRCR problem, we removed all tasks that appeared to have no feasible solution, and for SRLG-disjoint DCR problem, we kept only trap instances. After that, the total number of DRCR instances is 8498, the total number of trap instances for ”random” SRLG is 3404, and for ”star” SRLG is 3264. We consider non-trap instances trivial and thus not representative for evaluating the BTCS algorithm.
We show information on average number of links in each part of dataset in Tables I, II. For the real-world networks we also include information of their planarity according to the paper [42] as it can be important for further analysis.
@@ IV-B DRCR
All instances of the DRCR problem in the dataset from [22] are relatively easy to solve for the Pulse+ algorithm. So, we do not split it into parts by topology types but generate statistics for the entire dataset altogether. This dataset has a total number of 11245 tasks, and the results are presented in Table III.
@@
We run four algorithms on this dataset. The first algorithm in the table is the Pulse+, implemented by the authors of [22] and obtained on Github [30]. Second is the Pulse+, which we developed independently. Surprisingly, these two significantly differ in performance, which we assume to be due to implementation difference, since all main features of our Pulse+ implementation follow the ideas of [22]. Our code is built using Boost Graph Library, while the code in [30] uses its own data structure for graph representation.
@@ IV-C SRLG-disjoint DCR problem
To highlight algorithm performance when dealing with different network topologies and different SRLG assignment options, we present the results separately for real-world, ER, and SF networks, each with random and star SRLG generators. There are six combinations, and the results are listed in tables V - IX. The only combination missing in these tables is real-world networks with random SRLG, which has no feasible solutions. To show the advantage of parallel execution of BTCS and give further insight into its nature, we run it in 1, 2 and 4 parallel threads. For the corridor width in all experiments we use the cost of the shortest edge in the graph, multiplied by 10.
@@ IV-D Analysis and discussion
IV-D1 Single-thread performance
All results in the previous section are presented for single- and multithread implementation of the BTCS algorithm. Importantly, considering the number of feasible solutions found under the limit of 20 or 50 ms, BTCS algorithm in most cases has advantage over CoSE-Pulse+ even in single-thread version. The only exception is ER instances with random SRLG, where the single-thread performance is inferior (though still comparable). Considering other metrics, such as maximal, average and median time to feasible solution, BTCS in single thread also has comparable performance to CoSE-Pulse+, and outperforms it in certain scenarios (see Tables V - IX for details). This observation underpins the claim that the competitiveness of BTCS originates from its basic principles and not from any sophisticated multithread implementation. The only exception is the SF instances with star SRLG, where the maximal execution time of BTCS in single thread is almost 6x worse than that of CoSE-Pulse+.
IV-D2 Multithread performance
The multithread execution results in the previous section are obtained with a simple and relatively straightforward approach. In this implementation, the algorithm performs parallel search for AP candidates in several subsequent cost corridors. Other variants, such as the parallel search for disjoint PP within a corridor, are possible, as mentioned in Section III-D. Conflicting link paradigm also allows parallelization, as in [21], although it was not generalized to handle delay-constrained problems to the best of our knowledge. We consider detailed research of parallel execution capabilities for the SRLG-disjoint DRCR problem a subject of future work.
The approach implemented in the present work allows further insight into how the distribution of feasible solutions influences the algorithm’s performance. Note that Tables V - VII show ”normal” behavior, i. e. the number of found feasible solutions increases and execution time decreases with growing number of threads. On the contrary, Tables IX - IX show ”abnormal” behavior, namely, the performance deteriorates between two and four threads. We attribute this issue to the path distribution relative to cost. Too many threads means that the higher thread will often search in the space above the shortest feasible solution. Since there are more AP candidates at higher cost values, these threads will take more time, slowing down the whole process.
@@ IV-D2 Multithread performance
To understand this phenomenon more deeply, we introduce parameter $\alpha$ to account for the corridor width in the algorithm implementation. The width of the corridor is the cost of the shortest edge in the graph, multiplied by $\alpha$. So, the experiments performed in the previous section correspond to $\alpha=10$. In Figure 5, we plot the number of feasible paths found under 20 and 50 milliseconds as a function of $\alpha$, which is effectively the corridor width. The figure shows that algorithm performance in one and two threads is quite stable with respect to $\alpha$ and changes only a little in the wide range of parameter values. The value $\alpha=10$ is close to optimal for these two options, and double-thread implementation has clear advantage over single-thread. At the same time, the four-thread implementation has no overall advantage compared to double-thread, and its performance decays quickly with growing $\alpha$.
@@ IV-D3 Worst-case performance
Although worst-case performance is interesting theoretically, it is a less valuable metric in real applications. For an algorithm intended to be applied in the operation of modern networks, with acceptable delays measured in milliseconds, an algorithm that solves a problem in a second does not gain much advantage over the one that takes 5 seconds for the same task. This is why we decided to primarily compare algorithms by the number of problems solved in 20 and 50 milliseconds. Nonetheless, we provide and discuss some data below to give insight into the algorithm’s worst-case performance.
@@
Figure 6: Distribution of solution times for SRLG-disjoint DCR problem on the real-world network instances. (a): absolute execution time of BTCS algorithm on Zoo dataset with star SRLG. (b): distribution of execution times for CoSE-Pulse+ algorithm against single-thread BTCS. (c): BTCS algorithm execution time dependency on the graph size. (d): BTCS algorithm execution time dependency on the number of cost corridors explored before the solution is found. (e): BTCS algorithm execution time dependency on the number of AP candidates that are shorter than the solution, but have no disjoint protection.
@@ V Conclusion and future work
We suggest several ways of future improvement. One of them is application of the technique inspired by the queuing procedure from paper [15]. The original procedure stops pulses that have reached certain number of links, placing them in a queue and reusing again after all pulses have reached the limit. This approach turns the DFS search in Pulse algorithm into a combination of DFS and BFS, making it pure BFS when the depth limit is set to 1. This can be used as an inspiration in the search for AP candidates in BTCS, by placing the pulses that have reached the upper bound of current cost corridor in a queue and reusing in the next corridor instead of pruning them. It will improve speed at the price of memory to store all the pulses, that otherwise would be pruned.
Also it may be possible to use observations on search space structure made in this work to speed up the CoSE algorithm, or algorithms, developed for planar networks in [28] and [29]. A merging of these techniques, exploiting conflicting link sets or leveraging planar topology together with judiciously chosen cost bounds for efficient pruning, may result in further progress in the area.
Another way of improvement, considering the multithread implementation of the algorithm, is the development of more sophisticated ways of interaction between threads that will allow to stop execution when the solution is found in any thread. It will also allow mitigation of the performance decline for higher number of threads, observed on SF networks. Carefully designing breaking rules for multithread execution, one can also turn the algorithm into an approximate one, trading exactness of the solution for speed. For example, it can be achieved by independent exploration of odd and even corridors in two parallel threads, allowing one thread not to wait for the other on each step and breaking at the time when the first protected pair is found. Since the exploration goes in the bottom-top direction, the obtained solution will be relatively close to the optimal. We also can think of including parallelism on different level, and instead of searching for candidates in parallel corridors, search for protection paths in parallel for all AP candidates in each corridor, as was discussed in Section III-B.
Finally, in the process of this work we bumped into the question of having a reliable standard and realistic dataset to evaluate algorithms for modern communication standards. As we showed in the Experiment section, all available real-world networks are relatively small and easy to solve. The Topology Zoo collection was first published in 2011, and hardly reflects modern state of the networking technology. It also lacks of important data, such as real costs, delays and SRLG distributions. The Topology Zoo has been analyzed in the work [42] and the majority of networks were found to be planar. The same work observed the tendency that bigger networks are more often non-planar. The question remains, is it actually the case with real networks containing thousands of nodes?
In our work we generated costs, delays and SRLG, following the rules proposed in [22]. Although we tried our best to test different network topologies and SRLG patterns, it is still possible that the path distribution with cost on more realistic networks will not benefit our algorithm to the same degree as the generic dataset used in this work. Nonetheless, we still believe that any pruning-based algorithm can make use of the analysis of such distribution, as we did in this work for the first time.
@@ A-A Pulse+ algorithm
The pseudocode, presented here, corresponds to the original Pulse+ algorithm, proposed in paper [22]. It is intended to solve Delay-Range Constrained Routing problem and return the optimal path for it. In our paper it is used twice: in section III-B (Algorithms 3 and 4) and in Algorithm 5. We divide the code into two parts for convenience: the high-level code (Algorithm 1) and details of recursive Pulse function (Algorithm 2).
@@ 5      $current\_min\_cost=c$
Lines 2 and 2 of algorithm 2 contain infeasibility and optimality prunings, respectively. Dominance pruning is not applicable to the scenario with nonzero $d_{low}$, so, generally this algorithm is not as efficient as original Pulse. In order to avoid loops in found paths, it is also essential to track visited nodes in this scenario.
@@ A-B Pulse+ modification for all paths search in cost corridor
On the line 5 of algorithm 5 we use modification of Pulse+ algorithm to save all paths in current cost corridor. We put it into Appendix, because the difference with original Pulse+ is minor.
@@ A-C Pulse+ algorithm for first feasible path
In this section we describe, perhaps, the simplest implementation of the Pulse+ algorithm, used in algorithm 5 on line 5. It is intended to return any feasible path as fast as possible. So, it does not keep track of cost value, and applies only infeasibility pruning. It also must terminate the recursion, when the first feasible path was found. The function on line 11 of algorithm 11 is responsible for it, but we do not go into details of implementation. On practice it can be done by inclusion of a global boolean flag or by any other mean.